OpenStudy (anonymous):

An urn contains seven red balls, seven white balls, and seven blue balls, and sample of four balls is drawn at random without replacement. Compute the probability that all of the balls in the sample are the same color.

OpenStudy (perl):

Probability ( 4 balls of same color) = P ( RRRR or WWWW or BBBB ) =P(RRRR) + P(WWWW) + P(BBBB) can you calculate those separately

OpenStudy (perl):

P(RRRR) = 7/21 * 6/20 * 5/19*4/18

OpenStudy (perl):

same goes for P(WWWW) , and P(BBBB)

OpenStudy (perl):

R = red ball W = white ball B = blue ball

OpenStudy (anonymous):

is it 0.0174 final answer. for individual is 0.0058

OpenStudy (anonymous):

i tried it but it is wrong

OpenStudy (perl):

does it say to estimate

OpenStudy (perl):

the exact value is 1/57

OpenStudy (perl):

3* 7/21 * 6/20 * 5/19*4/18 = 1/57

OpenStudy (anonymous):

thank you perl

OpenStudy (anonymous):

its correct now